OpenStudy (freemap):

Heath wants 10 L of a solution that is 90% bleach. He will combine pure bleach with a mixture that is 20% bleach. (a) Write a system of equations to model the problem. Use p for the number of liters of pure bleach and m for the number of liters of the mixture that is 20% bleach. (b) Solve the system. Show your work. How many liters of pure bleach will Heath use? How many liters of the 20% mixture will he use?

OpenStudy (freemap):

@ganeshie8

OpenStudy (freemap):

@KamiBug

OpenStudy (freemap):

how will I write this?

OpenStudy (freemap):

@KamiBug

OpenStudy (freemap):

@welshfella

OpenStudy (freemap):

I need part A and B @welshfella

OpenStudy (welshfella):

2 equations p + m= 10 ( as total volume = 10 L) p + 0.20m = 9 ( amounts of pure bleach)

OpenStudy (freemap):

okay I'm going to solve that

OpenStudy (freemap):

okay I tried but I'm confused of how to solve that

OpenStudy (welshfella):

the easiest way is to subtract the second equation from the first This will eliminate p

OpenStudy (welshfella):

this will give 0.8m = 1

OpenStudy (freemap):

p+0.2m=9 p+m=10 cancel out p. 0.2m= 9-10=1 0.2m=1

OpenStudy (freemap):

how did you get 0.8?

OpenStudy (welshfella):

m - 0.2m = 0.8m

OpenStudy (welshfella):

and 10 - 9 = 1

OpenStudy (freemap):

okay I get 10-9=1 I'm sorry I just don't get it m-0.2m=0.8? If m was 1 it would be 1-0.2=.8 okay I get sorry

OpenStudy (freemap):

I didn't feel m as 1. I get it

OpenStudy (welshfella):

so m = 1 / 0.8 L then to find p subtract the result from 10

OpenStudy (freemap):

1.25-10=-8.75

OpenStudy (freemap):

@welshfella

OpenStudy (welshfella):

thats correct 8.75 L of pure bleach 1,25 L of 20% bleach

OpenStudy (welshfella):

* 1.25 L

OpenStudy (freemap):

Thanks @welshfella

OpenStudy (freemap):

1.25 L is How many liters of the 20% mixture will he use